# E-scooter effects on public transport demand: a case study in Santiago, Chile

### v1

@@ Abstract
As cities adopt sustainable mobility solutions, electric scooters (e-scooters) offer both challenges and opportunities for public transportation systems. This study, the first in Latin America, examines the effects of e-scooters on public transport demand in Santiago, Chile, focusing on two scenarios: “generation” of trips (trips starting in study zones) and “attraction” of trips (trips ending in study zones). A negative binomial regression model was applied to data from public transport smart cards and e-scooter GPS. The methodology included urban area clustering and a differences-in-differences approach. The findings reveal significant regional differences: in the Central Region, public transport trips decreased by 21.38% in the generation scenario, while bus trips increased by 76.39%. In the Intermediate Region, metro trips increased by 70.05%, and in the Peripheral Region, bus trips increased by 84.64%. These results suggest that e-scooters reduce public transport usage in highly accessible areas but increase it in less accessible regions.
@@
Electric scooter , e-scooters , Micro-mobility , Mobility as a Service , Public Transport , Public Transit
journal: Transportation Research Part D
@@
[aff1]organization=Universidad del Desarrollo, addressline=Av. Plaza 680, Las Condes, city=Santiago, country=Chile
@@ 1 Introduction
Micromobility refers to modes of transportation utilizing small, lightweight vehicles powered by human effort or electricity, operated at low speeds (Dediu, 2019). These vehicles include bike-sharing systems, both docked and dockless e-scooters, as well as other vehicles weighing less than 500 kg, designed for individual travel in urban environments. Micro-mobility works well in dense areas (Tuli et al., 2021) and can help reduce traffic, cut emissions, and improve local economies.
Among these micro-mobility solutions, shared e-scooters have emerged as a prominent option. These vehicles typically operate through a Mobility as a Service (MaaS) application (Narayanan and Antoniou, 2023), allowing people to use their mobile phones to locate nearby scooters, reserve one for their trip, and ride to their destination. Depending on the service, e-scooters can sometimes be left parked anywhere, while in other cases they must be parked in specific designated spots. These services are usually led by private companies rather than transportation authorities and have been marketed as complements or enablers of public transportation systems, primarily as a first- or last-mile transportation mode (Zuniga-Garcia et al., 2022).
Nevertheless, the full extent of e-scooters’ impact on urban environments remains insufficiently understood. This is due to their recent introduction, the frequent lack of adequate regulation, and the limited availability of data (Aarhaug et al., 2023). On one hand, they offer the potential for modal shifts, complementing public transportation systems  (Latinopoulos et al., 2021; Zuniga-Garcia et al., 2022) and providing first- and last-mile solutions. On the other hand, they present challenges such as the displacement of unused vehicles, unauthorized use of public spaces, safety concerns for both riders and pedestrians (Stigson et al., 2021), conflicts with pedestrians (Gössling, 2020), and environmental issues related to their disposal (Echeverría-Su et al., 2023).
The scarcity of comprehensive studies on the overall impact of e-scooters complicates our understanding of their effects on urban environments, especially in relation to other transportation modes. Most existing research on this topic focuses on cities in the United States and Europe, which may not fully capture the complex interactions between e-scooters and public transport systems in other regions. Notably, in South America, there is a significant research gap in this field (Mitropoulos et al., 2023). This gap is particularly critical in cities like Santiago, where recent studies have shown significant changes in transportation patterns, including a reduction in mass-transit usage and an increase in motorized transport and taxi usage (Graells-Garrido et al., 2023a).
In this context, our study aims to address this critical gap by examining the effect of MaaS e-scooter availability on public transport usage in Santiago, Chile. Using a differences-in-differences methodology, we analyze official public transport data alongside e-scooter trip information from 2019, the peak usage period for these vehicles in Santiago. This is the first study of its kind in Latin America, offering empirical evidence on the interaction between e-scooters and public transport in a regional context. Our analysis provides valuable insights for urban planners and policymakers seeking to develop more effective and sustainable urban mobility strategies. By focusing on both trip generation and attraction scenarios, we capture the varying effects of e-scooter usage across different urban regions. Additionally, we apply a novel clustering method to automatically determine treatment and control zones, ensuring robust comparisons across different urban contexts. Furthermore, this study expands the scope of prior research by examining a multimodal public transport network, rather than focusing on a single mode, providing a comprehensive understanding of the interactions between e-scooters and public transport.
The paper structure is as follows: the next section presents a summary of the existing literature on the impact of shared e-scooters on public transportation usage. Following this, the study context and data description are presented. The methodology section is then detailed, divided into the study variables, clustering of zones of study, and the application of the differences-in-differences approach for the generation and attraction travel scenarios. This is followed by the results and discussion for these scenarios. Next, the paper addresses limitations and explores areas for improvement and future research directions. Finally, the paper concludes with a summary of the findings and their implications for urban mobility strategies
@@ 2 Literature Review
Most studies using e-scooter trips rely on spatial analysis and perform spatial and statistical analyses of trip data to estimate the relationship between e-scooter and public transportation trips (Vinagre Díaz et al., 2023; Yan et al., 2021; Aarhaug et al., 2023; Bai and Jiao, 2020; Jiao and Bai, 2020; Hosseinzadeh et al., 2021). Conversely, few studies have employed a quasi-experimental setting to understand this relationship (Ziedan et al., 2021a, b; Luo et al., 2021; Tyndall, 2022; Weschke et al., 2022).
@@
Several spatial analysis studies have classified e-scooter trips as either complementary or substitutive to the transit system based on the proximity of e-scooter trip origins and destinations to transit stations (Vinagre Díaz et al., 2023; Yan et al., 2021; Aarhaug et al., 2023; Bai and Jiao, 2020; Hosseinzadeh et al., 2021; Jiao and Bai, 2020). For instance, Vinagre Díaz et al. (2023), based on data from Rome, reported that 50% of e-scooter trips either replace public transport or serve areas with inadequate public transport infrastructure. Similarly, Yan et al. (2021), using data from Washington, DC, found that approximately 70% of e-scooter trips could be considered substitutes, while around 20% were complementary to the transit system. Aarhaug et al. (2023) highlighted that e-scooters can simultaneously act as competitors and complements to public transport, based on their research conducted in Norwegian cities. Bai and Jiao (2020) utilized regression-based models to show that e-scooters can serve both roles in public transport systems in Austin and Minneapolis. Additionally, Hosseinzadeh et al. (2021) demonstrated a positive correlation between e-scooter usage and the density of bus stops in Louisville, indicating that e-scooters can complement public transport infrastructure, while Jiao and Bai (2020) confirmed that proximity to public transport stations is linked to increased e-scooter usage in Austin.
Regarding quasi-experimental analysis, there are two groups of studies. One group focuses on the effect of scooter trips on the number of bus riders  (Luo et al., 2021; Ziedan et al., 2021a, b), while the second group evaluates the demand for e-scooter trips based on light rail systems  (Tyndall, 2022; Weschke, 2023). Luo et al. (2021) conducted a spatial-temporal analysis using difference-in-difference models to assess the impacts of shared e-scooters on bus usage in Indianapolis, finding that 27% of e-scooter trips potentially competed with buses, while 29% were likely to increase bus usage. Ziedan et al. (2021a) used fixed-effects regression to investigate the effects on bus usage in Louisville, noting that shared e-scooters did not significantly impact local bus usage. Conversely, Ziedan et al. (2021b) classified trips by trip purpose using data from Nashville, Tennessee. The study found that routine weekday trips are related to a 0.94% decrease in bus usage, while social e-scooter trips, are related to a 0.86% increase in weekday bus usage.
Tyndall (2022) analyzes e-scooter trip data in Seattle, while Weschke (2023) uses data from Berlin. Tyndall (2022) explores the expansion of a light rail system in the U.S. as a natural experiment, applying a difference-in-difference approach. The study finds that the introduction of a new light rail station led to a significant increase in the presence of shared bicycles and e-scooters within a five-minute walking distance from the station. Similarly, Weschke (2023) examines the extension of a metro line in Berlin, Germany, through a quasi-experiment, revealing that e-scooter demand increases at newly opened stations, while it significantly decreases when a station is closed. Both studies concluded that public transport acts as a complement for shared e-scooters.
Table 1 below provides a summary of the key studies that examine the relationship between e-scooter trips and public transportation, highlighting their methodologies, data, and main findings.
Authors Method Data Place Main result
Vinagre Díaz et al. (2023) Clustering technique E-scooter trips, and rail-based transit network Rome, Italy 50% of scooter trips either replace transit or cover areas with low transit accessibility
Yan et al. (2021) Spatiotemporal analysis E-scooter trips, bikeshare network, and transit network Washington, DC, US 20% of trips are last-mile connections to transit and 70% are transit substitutes
Ziedan et al. (2021a) Fixed effects regression models E-scooter trips, network and trips bus ridership data Louisville, US Complementary effect on express bus route - transit increase 5 trips for every 10 e-scooter trips
Luo et al. (2021) Spatial-temporal analysis with a Difference-in-Differences model E-scooter trips, network and trips bus ridership data Indianapolis, Indiana, US 27% of e-scooter trips compete with bus trips and 29% complement the bus system
Ziedan et al. (2021b) Fixed effects regression models E-scooter trips, and trips bus ridership data Nashville, Tennessee, US Net impact of e-scooter trips is -0.08% on bus ridership
Aarhaug et al. (2023) Logit regression model E-scooter trips, travel survey, planning transport data Oslo, Norway 50.5% of e-scooter trips compete with public transport and 20% complement public transport
Baek et al. (2021) Mixed logit model Stated Preference survey Seoul, South Korea 40.6% of trips indicating a preference for e-scooters, 39.2% for buses, and 20.2% for walking
Bai and Jiao (2020) Spatial analysis and negative binomial regression models E-scooter trips, build environment, and demographic data Austin, Texas, US; Minneapolis, Minnesota, US 52% and 47% (Austin and Minneapolis) of e-scooter trips compete with public transport and 31% and 29% of them complement public transport
Jiao and Bai (2020) Spatial analysis and a negative binomial regression model E-scooter trips, build environment, and demographic data Austin, Texas, US 52% of e-scooter trips compete with public transport and 31% of them complement public transport
Hosseinzadeh et al. (2021) Geographically Weighted Regression E-scooter trips, build environment, and demographic data Louisville, Kentucky, US 50% of e-scooter trips substituting public transportation and 25% complementing it
Tyndall (2022) Difference-in-difference E-scooter and e-bikes parking locations, transit network, demographic, and geographic data Seattle, Washington, US The flow of new micromobility vehicles increased significantly within a 5 min walking radius of the new rail station
Weschke (2023) Difference-in-differences E-scooter trips, OpenStreetMap Data Berlin, Germany Public transport acts as complement for shared e-scooters
Table 1: Summary of Studies on the Impact of E-Scooter Trips on Public Transportation. This table summarizes studies on the impact of e-scooter trips on public transportation. The columns include the authors and year of publication, the methodological approach (e.g., clustering, spatiotemporal analysis, regression models), the types of data used (e.g., e-scooter trips, network data, bus ridership, travel surveys), the geographic location of the study, and the main findings regarding how e-scooter trips affect public transportation.
These previous studies have laid the foundation for understanding the interaction between e-scooters and public transport. Building on their insights, our study expands prior research (Luo et al., 2021; Ziedan et al., 2021a, b; Tyndall, 2022; Weschke, 2023) by employing a quasi-experimental approach. The key differences and contributions of our study are:
  • 1.
    Our study is conducted in Santiago, Chile, offering a Latin American perspective that is underrepresented in the existing literature. This regional focus is crucial, given the unique urban dynamics and public transport challenges in Latin American cities.
  • 2.
    We consider a multimodal public transport network, whereas Ziedan et al. (2021a, b), Luo et al. (2021), Tyndall (2022), and Weschke (2023) focus solely on specific transit modes.
  • 3.
    Our study encompasses a large-scale transport network, covering the entire city of Santiago, including central, intermediate, and peripheral regions. This comprehensive approach allows us to capture a wider range of interactions between e-scooters and public transport across different urban settings.
  • 4.
    We propose a geographical and transport of mode analysis to gain a deeper understanding of which public transport trips are related to e-scooter trips.
  • 5.
    Our study proposes a clustering methodology that automatically determines control and treatment zones, as opposed to other studies that select these zones manually.
@@ 3 Data Description
This section describes the data sources used in our study of transportation demand in the Great Metropolitan Area of Santiago, Chile. We analyze both public transport and shared e-scooter usage across two distinct time periods, utilizing comprehensive datasets to examine the impact of e-scooters on public transport demand.
Santiago, Chile’s capital and largest city, had a population of nearly 8 million and covered an urban area of 867.75 square kilometers as of the 2017 Census. The city comprises over 40 independent municipalities, with 35 in the urban core and the rest on the periphery (see Fig. 1 a). Santiago’s public transportation system is integrated and multimodal, featuring a nearly flat fare structure across metro, urban buses, and a rail service. This system allows up to three transfers within a two-hour window. The bip! smart card is the sole payment method accepted, requiring validation when boarding buses or entering metro stations, but not upon exiting. Consequently, the smart card data only records trip origins. In addition to public transport, Santiago offers extensive taxi services, several ride-hailing applications, and a public bike-sharing system.
For this work, we analyze two types of transportation demand: public transport and shared e-scooters. Our analysis focuses on two distinct time periods:
  1. 1.
    Public transport data: We examined bip! records from the second week of May 2018 and the second week of May 2019. These weeks were chosen as they are similar in terms of typical urban activity, without any particular events or interventions in the city. It is important to note that by 2019, the newly inaugurated Metro Line 3 was already in operation, which we considered in our analysis.
  2. 2.
    E-scooter data: We analyzed trips from a shared e-scooter application, covering the period from October 10th, 2018 to April 15th, 2019. This dataset includes the period when e-scooter services were introduced and became established in Santiago.
We selected these specific time periods to capture both the typical urban activity patterns and the introduction and establishment of e-scooter services in Santiago. The inclusion of data from 2018 and 2019 also allows us to account for the impact of the newly inaugurated Metro Line 3, which began operations between these two periods.
In this study, the unit of analysis is the geographical zoning of the Metropolitan Public Transport Directory (DTPM). It consists of 804 transportation analysis zones (see Fig. 1 b). These zones have an average area of 1.23 Km2. Next, we describe the datasets used to measure the effect of e-scooters in public transport demand.
@@
Our primary dataset originates from the ADATRAP system (Gschwender et al., 2016), provided by DTPM. ADATRAP consolidates multiple data sources from Santiago’s public transport system, including smart card transactions, bus GPS records, and vehicle operator information, among others. These data are processed and enhanced by analyzing user travel patterns and the spatial distribution of distances to boarding points (Munizaga and Palma, 2012). Trips are reported at bus stop or metro station level, which can be aggregated into zones (see Fig. 2.a to view the spatial distribution of all trips under study), municipalities or macro-areas that contain multiple municipalities.
Figure 2: Distributions of Public Transport trips. Left (a): a choropleth map of total trips with respect to Traffic Analysis Zones. Top right (b): total trips per day, per year. Bottom right (c): relative distribution of trip starting periods for each day, where the periods are defined in Table A.6.
Through an agreement between a e-scooter operator and the Las Condes municipality in Santiago (see Fig. 4), we were able to analyze the origins and destinations of e-scooter trips in Las Condes. The e-scooter operator provided a dataset of GPS traces. Trips in this dataset were detected using a stop-based algorithm (Graells-Garrido and Saez-Trumper, 2016). Each observation indicated whether the vehicle was being under control of a traveler or a staff person, as such, we were able to discard relocation trips. In total, this dataset contains 45K trips between October 10th, 2018, and April 15th, 2019. The time of the day and distance distributions of those trips (and ADATRAP trips for comparison) are shown in Fig. 3. Notably, e-scooter trips have a less pronounced morning peak (6:00-9:00 hrs) in comparison to public transport trips, although the afternoon peak (5:30-20:30 hrs) shows a similar trend. At lunch time (12:00-14:00 hrs), e-scooter trips present a more pronounced distribution than public transport, which is arguably related to the area of availability, as Las Condes is a municipality that concentrates high-density working areas and business centers. Notably, trips in e-scooters are shorter than public transportation trips (with means of 1.44 Km and 9.25 Km, respectively). The origins and destinations of trips seem to be closely related to the metro stations available in their area of operation (see Fig. 4).
@@
Figure 4: Origins and destinations of shared e-scooter trips. The top row shows trip origins in three periods of the day (morning peak, lunch and afternoon peak), whereas the bottom row shows trip destinations for the same period. The map on the right highlights the area of availability of the e-scooter service: the municipality of Las Condes.
Finally, we utilized demographic data from the 2017 Chilean Census  (National Institute of Statistics, 2017). The variables analyzed included population counts by age and sex, as well as the average educational level (measured in years of schooling). Educational level was used as a proxy for income to capture the socio-economic characteristics of each zone. As the census does not report individual income directly, we used years of schooling as a proxy for income. This approach is supported by research showing strong correlations between educational attainment, income distribution, and inequality (Gregorio and Lee, 2002; Jerrim and Macmillan, 2015). All variables were aggregated at the traffic analysis zone level. We have also used the official General Transit Feed Specification (GTFS) provided by DTPM to calculate some accessibility indicators. We describe these datasets in the next section, as their variables are intricately related to our methodology to quantify the effect of shared e-scooters on public transport demand.

### v2

# Impact of Shared E-scooter Introduction on Public Transport Demand: A Case Study in Santiago, Chile

@@ Abstract
This study examines how the introduction of shared electric scooters (e-scooters) affects public transport demand in Santiago, Chile, analyzing whether they complement or substitute for existing transit services. We used smart card data from the integrated public transport system of Santiago and GPS traces from e-scooter trips during the initial deployment period. We employed a difference-in-differences approach with negative binomial regression models across three urban regions identified through k-means clustering: Central, Intermediate, and Peripheral. Results reveal spatially heterogeneous effects on public transport boardings and alightings. In the Central Region, e-scooter introduction was associated with significant substitution effects, showing a 23.87% reduction in combined bus and metro boardings, suggesting e-scooters replace short public transport trips in high-density areas. The Intermediate Region showed strong complementary effects, with a 33.64% increase in public transport boardings and 4.08% increase in alightings, indicating e-scooters successfully serve as first/last-mile connectors that enhance transit accessibility. The Peripheral Region exhibited no significant effects. Metro services experienced stronger impacts than bus services, with metro boardings increasing 9.77% in the Intermediate Region. Our findings advance understanding of micromobility-transit interactions by demonstrating that both substitution and complementarity can coexist within the same urban system, depending on local accessibility conditions. These results highlight the need for spatially differentiated mobility policies that recognize e-scooters’ variable roles across urban environments.
@@
Electric scooters , e-scooters , Micromobility , Mobility as a Service , Public Transport , Public Transit
journal: Travel Behaviour and Society
@@
[aff1]organization=Universidad Técnica Federico Santa María, addressline=Avenida Vicuña Mackenna 3939, San Joaquín, city=Santiago, country=Chile
@@ 1 Introduction
Cities worldwide need to understand how shared electric scooters (e-scooters) affect public transport as they invest in sustainable mobility solutions. In dense urban areas, micromobility (the use of small, lightweight vehicles weighing less than 500 kg (Dediu, 2019)) can reduce traffic, cut emissions, and improve local economies (Tuli et al., 2021). However, the impact of e-scooters on transit ridership remains unclear. Previous studies show mixed effects: some report reduced public transport demand (Vinagre Díaz et al., 2023), while others find e-scooters acting simultaneously as competitors and complements to transit (Yan et al., 2021; Aarhaug et al., 2023).
Shared e-scooters have emerged as a prominent micromobility option through Mobility as a Service (MaaS) applications (Narayanan and Antoniou, 2023). These services let users locate, reserve and ride scooters via mobile phones, with flexible parking options. Private companies typically operate these services and market them as complementary to public transit, particularly for first/last mile connections (Zuniga-Garcia et al., 2022). While they could enable modal shifts and complement public transit (Latinopoulos et al., 2021; Zuniga-Garcia et al., 2022), they also raise concerns about vehicle displacement, unauthorized space use, safety (Stigson et al., 2021; Graells-Garrido et al., 2025), pedestrian conflicts (Gössling, 2020), and environmental impacts (Echeverría-Su et al., 2023).
Current knowledge gaps are significant. Most research analyzes e-scooter effects in specific city areas or transit lines, but no study examines how effects vary across different urban structures such as central business districts versus peripheral zones. Additionally, most studies focus on single public transport modes (either bus or metro) rather than evaluating effects on integrated multimodal systems. This represents an important limitation because e-scooters may impact central areas differently than suburban areas, and may affect bus services differently than metro services.
Our study examines three critical research questions: How does introducing shared e-scooters affect public transport ridership across different urban contexts? Do shared e-scooters effectively solve first/last mile challenges for public transport? How do these effects vary by transport mode (bus vs. metro) and urban region? The answers can help urban planners and transit agencies make informed decisions about e-scooter regulations, infrastructure planning, and integration with existing transit systems.
Santiago, Chile provides an ideal case study. The city has seen significant transport changes, including reduced transit use and increased motorized transport (Graells-Garrido et al., 2023a), while introducing e-scooters during our study period. Most research examines U.S. and European cities, leaving gaps in understanding how e-scooters affect transit in other contexts. This gap is especially relevant in South America (Mitropoulos et al., 2023), where cities like Santiago face different urban dynamics and transit challenges than well-studied North American and European contexts.
Using a difference-in-differences methodology with comprehensive smart card and GPS data, we advance the field in three ways. First, we demonstrate that both substitution and complementarity coexist spatially within the same urban system, with substitution dominating in transit-rich central areas (-23.9% in boardings) and complementarity emerging in intermediate zones (+33.6% in boardings). This spatial heterogeneity explains why previous studies have reached different conclusions across various urban contexts. Second, we contribute methodologically by developing an automated clustering approach using k-means regionalization that eliminates arbitrary manual selection of treatment and control zones common in prior quasi-experimental studies, enhancing the validity of causal inference. Third, we provide the first systematic evidence of mode-specific effects (bus vs. metro) in a Latin American context, revealing that metro services show stronger complementary relationships with e-scooters than bus services.
The paper proceeds as follows: Section 2 reviews literature on e-scooter impacts. Section 3 describes our data and study context. Section 4 details our methodology, including clustering and regression approaches. Section 5 presents results, followed by discussion in Section 6. Section 7 concludes with implications for urban mobility strategies.
@@ 2 Literature Review
Research on e-scooter impacts employs two primary methodological approaches: spatial analysis and quasi-experimental methods. Each approach reveals different aspects of the e-scooter-transit relationship, but significant gaps remain in understanding spatial heterogeneity and multimodal effects.
Spatial Analysis Studies
@@
Spatial analysis studies classify e-scooter trips as complementary or substitutive based on proximity to transit stations. These studies consistently find mixed effects within the same urban areas. Research in Rome found 50% of e-scooter trips either replace public transport or serve areas with inadequate transit infrastructure (Vinagre Díaz et al., 2023). Analysis in Washington, DC showed approximately 70% of trips substituting for transit while 20% complemented it (Yan et al., 2021). Studies in Norwegian cities demonstrated that e-scooters simultaneously compete and complement transit (Aarhaug et al., 2023). Research in Austin and Minneapolis found 47–52% of trips competing with transit and 29–31% complementing it (Bai and Jiao, 2020). Additional studies in Austin showed increased e-scooter use near transit stations (Jiao and Bai, 2020), while Louisville research found positive correlation between e-scooter usage and bus stop density, suggesting complementary effects (Hosseinzadeh et al., 2021). A different methodological approach using stated preference surveys in Seoul found that 40.6% of travelers preferred e-scooters, 39.2% chose buses, and 20.2% opted for walking (Baek et al., 2021). Additional studies in Austin showed increased e-scooter use near transit stations (Jiao and Bai, 2020).
These mixed findings suggest that spatial context mediates e-scooter-transit relationships, yet most studies aggregate effects across entire urban areas without examining how local characteristics influence outcomes.
Quasi-Experimental Studies
Quasi-experimental research provides stronger causal evidence but focuses narrowly on specific modes or routes. Bus-focused studies show varying results: research in Indianapolis found 27% of e-scooter trips competed with buses while 29% increased bus usage (Luo et al., 2021). Studies in Louisville showed complementary effects on express routes (Ziedan et al., 2021a), while analysis in Nashville found different effects by trip purpose (Ziedan et al., 2021b).
Rail-focused studies consistently show complementary effects. Research in Seattle found new rail stations significantly increased nearby e-scooter presence (Tyndall, 2022). Studies in Berlin showed e-scooter demand rising near new stations and falling near closed ones (Weschke, 2023).
Research Gaps
Current literature exhibits three critical limitations. First, studies examine either spatial patterns or causal effects, but not both within a comprehensive framework that accounts for urban heterogeneity. Second, most research focuses on single transport modes, missing interactions within integrated multimodal networks. Third, research concentrates on North American and European cities, leaving substantial gaps in understanding different urban contexts, particularly in Latin America where transit systems and urban development patterns differ significantly. Our study addresses these gaps by combining quasi-experimental methodology with systematic spatial analysis across an integrated multimodal network in Santiago, Chile. Unlike previous work that manually selects treatment and control zones, we employ automated clustering to enhance methodological rigor while examining both bus and metro interactions across different urban contexts within a single analytical framework.
@@ 3 Data Description
This study examines how e-scooters affect public transport demand in Santiago, Chile’s capital and largest city. As of the 2017 Census, Santiago had a population of nearly 8 million across 867.75 square kilometers, comprising over 40 independent municipalities, with 35 in the urban core and the rest on the periphery (see Fig. 1 a).
We analyze two primary data sources across specific time periods. First, we examined bip! smart card records from the second week of May 2018 and 2019, capturing typical urban activity without special events. The 2019 period includes the newly opened Metro Line 3. Second, we analyzed GPS traces from shared e-scooters between October 10th, 2018 and April 15th, 2019, covering the introduction and establishment of these services in Santiago. The dataset includes dockless e-scooters that users could park freely within service areas.
Our analysis uses geographical zones defined by the Metropolitan Public Transport Directory (DTPM), comprising 804 transportation analysis zones averaging 1.23 squared km each (see Fig. 1 b).
@@
Santiago’s public transport system integrates metro, urban buses, and rail with a nearly flat fare structure. The system allows up to three transfers within two hours using the bip! smart card (the only accepted payment method). The card records trip origins but not destinations, as validation occurs only when boarding buses or entering metro stations. Beyond public transport, the city offers taxi services, ride-hailing, and bike-sharing.
The public transport data comes from DTPM’s ADATRAP system (Gschwender et al., 2016), which consolidates smart card transactions, bus GPS records, and operator information. ADATRAP processes these data using travel pattern analysis and spatial distribution of boarding points (Munizaga and Palma, 2012). We can aggregate trips from bus stops or metro stations into zones (Fig. 2.a shows trip distribution), municipalities, or macro-areas.
Figure 2: Distributions of Public Transport trips. Left (a): a choropleth map of total trips with respect to Traffic Analysis Zones. Top right (b): total trips per day, per year. Bottom right (c): relative distribution of trip starting periods for each day, where the periods are defined in Table 8.
Through an agreement with Las Condes municipality (Fig. 4), we obtained GPS traces of 45K e-scooter trips. Using a stop-based algorithm (Graells-Garrido and Saez-Trumper, 2016), we identified trips and excluded staff relocation movements. The data reveals distinct patterns from public transport, as shown in Fig. 3. E-scooter usage shows less pronounced morning peaks from 6:00–9:00, while maintaining similar afternoon peaks between 17:30–20:30. We observed higher lunchtime usage between 12:00–14:00, likely due to Las Condes’ concentration of business areas. Shared e-scooter trips are notably shorter, averaging 1.44 km compared to 9.25 km for public transport, with origins and destinations clustering near metro stations.
@@
Figure 4: Origins and destinations of shared e-scooter trips. The top row shows trip origins in three periods of the day (morning peak, lunch and afternoon peak), whereas the bottom row shows trip destinations for the same period. The map on the right highlights the area of availability of the shared e-scooter service: the municipality of Las Condes.
We complemented these transport datasets with demographic data from the 2017 Chilean Census (National Institute of Statistics, 2017), including population counts by age and sex, as well as average educational level measured in years of schooling. We used educational level as a proxy for income, an approach supported by research linking education to income distribution (Gregorio and Lee, 2002; Jerrim and Macmillan, 2015). Additionally, we used DTPM’s General Transit Feed Specification (GTFS) data to calculate accessibility indicators, which we detail in the methodology section as they relate directly to measuring shared e-scooter effects on transport demand.